### v2

# SoK: Membership Inference Attacks on LLMs are Rushing Nowhere (and How to Fix It)

Abstract
Whether Large Language models (LLMs) memorize their training data and what this means, from privacy leakage to detecting copyright violations — has become a rapidly growing area of research over the last two years. In recent months, more than $10$ new methods have been proposed to perform Membership Inference Attacks (MIAs) against LLMs. Contrary to traditional MIAs which rely on fixed -but randomized- records or models, these methods are mostly evaluated on datasets collected post-hoc. Sets of members and non-members, used to evaluate the MIA, are constructed using informed guesses after the release of a model. This lack of randomization raises concerns of a distribution shift between members and non-members. In the first part of this paper, we review the literature on MIAs against LLMs. While most work focuses on sequence-level MIAs evaluated in post-hoc setups, we show that a range of target models, motivations and units of interest have been considered in the literature. We then quantify distribution shifts present in the $6$ datasets used in the literature, ranging from books to papers, using a bag of word classifier. Our analysis reveals that all of them suffer from severe distribution shifts. This challenges the validity of using such setups to measure LLM memorization and may undermine the benchmarking of recently proposed methods.
Yet, all hope might not be lost. In the second part of this work, we introduce important considerations to properly evaluate MIAs against LLMs and discuss potential ways forward: randomized test splits, injections of randomized (unique) sequences, randomized finetuning, and several post-hoc control methods. While each option comes with its advantages and limitations, we believe they collectively provide solid grounds to guide the development of MIA methods and study LLM memorization. We conclude by proposing and releasing comprehensive, easy-to-use benchmarks for sequence- and document-level MIAs against LLMs (see Table 6 and the datasets and code we release [^1] Link to the Github repository.. LLM memorization is an important and multi-faceted question, yet meaningful progress can only be achieved with the use of robust, independent benchmarks such as the ones we propose here.
1 Introduction
The prevalence of Large Language Models (LLMs) [77, 3, 43] has sparked significant interest in studying LLM memorization to better understand how they acquire their capabilities and learn facts. LLMs can also be trained on confidential or sensitive personal data, making the study of LLM memorization critical from both an information security and a privacy perspective [9, 71, 15, 65, 16, 72]. Similar research helps assess whether LLMs are being trained on copyrighted material [86, 100, 70, 28] and the extent to which they might generate such content [46]. Further, to fairly evaluate newly released models, it is crucial to understand whether they are trained on (and memorize) evaluation benchmarks [66, 56, 78, 33, 22, 54, 25].
Membership Inference Attacks (MIAs) have become a widely used tool to study the memorization of machine learning (ML) models [88, 12]. In an MIA, an attacker aims to infer whether a certain data sample has been used to train the target model. Recently, MIAs have also been proposed for LLMs [16, 106, 68, 86, 70, 109, 101, 62, 105, 108, 97, 73, 45], trying to infer whether a specific piece of text has been used to train an LLM. In most of this work, the unit of interest to infer membership has been a sequence (typically $\leq 256$ tokens), while other work has considered document [69, 86], to even dataset-level MIAs [67]. Beyond MIAs, prior work has also widely studied training data extraction from LLMs, either verbatim [13, 74, 112] or not [40, 80, 46], to measure LLM memorization.
Independent of how MIAs against ML models (e.g. image classification) are developed, they are traditionally evaluated in randomized experiments [88, 12], instantiated as a privacy game between a model developer and an attacker. Ye et al. [104] distinguishes between privacy games for fixed records and for fixed models. For fixed records, multiple models are trained with the target randomly included or excluded (shadow modeling) [63, 64, 91, 41]. Fixed model scenarios use a single model trained on randomly sampled, IID records [76, 84, 98]. While feasible for traditional ML models, applying a proper privacy game for LLMs presents unique challenges. Indeed, the mere scale and cost of training prohibits the development of LLMs solely for the purpose of developing proper privacy games. Further, it remains uncertain how findings from small-scale MIA experiments translate to the scale of real-world LLMs.
As a result, recent MIAs against LLMs have been evaluated in the absence of a randomized privacy game, relying instead on observational data collected post-hoc. Two concurrent works, published at ICLR [86] and USENIX Security [69], instantiate the membership inference task post-hoc; where an attacker aims to apply an MIA on a released LLM without access to the randomized training data distribution. Authors use publicly available datasets widely used for LLM training to collect likely members, and use data released on the same sources but after the LLM’s training data cutoff date to collect likely non-members. In this setup, both contributions report MIAs to work very well, reaching AUCs up to $0.8$. Further work has either used the exact same datasets, or similar datasets collected post-hoc, to develop and evaluate MIAs against LLMs [109, 101, 62, 105, 108, 97, 73, 45], all reaching high AUC values.
However, recently, strong concerns have been raised on the lack of randomization in post-hoc collection of members and non-members. Duan et al. [27] first suggested that this setup introduces a distribution shift between members and non-members due to the temporal gap in data collection, a concern later echoed by Maini et al. [67]. This shift would lead membership classifiers to learn temporal differences in language rather than membership, potentially inflating AUC scores and leading to an incorrect understanding of LLM memorization.
Contribution. In this work, we provide an overview of the recent surge in research on MIAs against LLMs and quantify any distribution shift in the evaluation setups used. We find that the absence of randomization has both a substantial impact on the measured MIA performance and that this issue is widespread. We then carefully examine the options towards more rigorous MIA evaluation to, ultimately, improve our understanding of LLM memorization.
First, we formalize the task of MIAs against LLMs, distinguishing between MIAs targeting LLM pretraining and finetuning, on the sequence-, document- and dataset-level. In Table 1, we review the literature, with a focus on the most recent advances (’23-’24) while including seminal initial work. MIA studies either merely evaluate MIAs from the literature on specific setups or also propose novel methods. We find that only in ’24, at least $10$ new MIA methods have been proposed. Next, we find that the primary motivation for such studies has evolved over time, with initial work mostly studying LLM finetuning for privacy and security reasons [15, 16, 71, 68, 65], while more recent work studies MIAs also for copyright and data contamination purposes, increasingly considering LLM pretraining. We then carefully examine which datasets have been considered to evaluate MIAs across the literature. We find that the majority of recent work does not consider randomized privacy games, and instead relies on data collected post-hoc (flagged with in Table 1).
Second, across datasets collected post-hoc, we examine the distribution shift between members and non-members as flagged by prior work [27, 67] to understand its impact on previously reported results. We propose to train a bag of words classifier to distinguish members from non-members, without considering the model [^2] Concurrently with this work, Das et al. [21] applies ‘blind’ baselines to datasets used for MIA evaluation. Authors confirm that simple classifiers can effectively distinguish members from non-members without taking the target model into account.. We apply this to data collected post-hoc [69, 86, 57, 105] and confirm that the simple classifier yields an AUC often better than proposed MIAs. This shows that there exists a serious distribution shift in language between the collected members and non-members - making it impossible to exclusively attribute the MIA performance to LLM memorization. This questions the effectiveness of newly proposed MIAs [109, 101, 62, 105, 108, 97, 73, 45] (latest one from September, 2024), potentially leading to false conclusions on LLM memorization.
Third, we discuss what a randomized setup to evaluate MIAs against LLMs entails. The cost of LLM development prohibits training multiple models for MIA evaluation, leaving only the privacy game for a fixed model as a practical option. Before examining the potential solutions, we elaborate on two core challenges. First, the inherent overlap between texts collected from the internet challenges the binary notion of membership. Prior work found significant overlap between randomized splits from the training data, either trying to achieve a clean model validation set [10, 99, 26, 95] or a useful set of members and non-members to evaluate MIAs [27]. Second, we elaborate on the unit of interest considered in MIAs against LLMs, ranging from a sequence to an entire dataset.
Lastly, we carefully examine the approaches available for rigorous evaluation of MIAs against LLMs in the future. We consider the following $4$ options and elaborate on their advantages and limitations:
  1. 1.
    The use of the randomized split of train and test data released for open-source LLMs as members and non-members for MIA evaluation, as suggested by prior work [55, 27, 67].
  2. 2.
    The injection of randomized unique sequences in LLM training data, as previously used to study LLM memorization [36, 15, 92, 93, 100, 70].
  3. 3.
    The creation of a randomized evaluation setup by finetuning LLMs on smaller datasets, as proposed before [68, 71, 107, 51, 65].
  4. 4.
    The application of control methods to make evaluation and calibration of MIAs against LLMs feasible on data collected post-hoc, either with a skewed prior [42, 47], de-biasing techniques [29] or techniques inspired by Regression Discontinuity Design (RDD) [50].
We conclude with an overview of recommended approaches to benchmark newly proposed MIAs in the future (Table 6). We hope this overview contributes to a better understanding of the recent advances in MIAs against LLMs and provides a foundation for future research to rigorously evaluate and interpret MIA results.
Target Unit of interest Motivation Dataset
MIA study
Finetuning
Pretraining
Sequence
Document or dataset
New MIA method
Privacy & Security
Copyright
Dataset contamination
WikiMIA* [86]
Arxiv*
Books*
Stack Exchange*
Pile [31]
MIMIR1 [27]
Randomized sequence
(’19) Carlini et al. [15]
(’21) Carlini et al. [16]
(’22) Mireshghallah et al. [71]
(’23) Mattern et al. [68]
(’23) Lukas et al. [65]
(’23) Shi et al. [86]
(’23) Meeus et al. [69]
(’23) Li et al. [55]
(’24) Duan et al. [27]
(’24) Zhang et al. [109]
(’24) Xie et al. [101]
(’24) Liu et al. [62]
(’24) Ye et al. [105]
(’24) Zhang et al. [108]
(’24) Wang et al. [97]
(’24) Mozaffari et al. [73] 2
(’24) Kaneko et al. [45]
(’24) Duarte et al. [28]
(’24) Miani et al. [67]
(’24) Wei et al. [100] 3
(’24) Meeus et al. [70]
(’24) Panaitescu et al. [79]
    Red indicates likely presence of the distribution shift
  • 1
    While MIMIR originates from a randomized setup, subsquent deduplication [27] could introduce a distribution shift (Sec. 6.1).
  • 2
    Custom datasets with the same underlying data as the indicated dataset.
  • 3
    SHA-hashes extracted from the StackExchange datasets, likely doesn’t suffer from the distribution shift.
Table 1: Overview of prior work studying MIAs against LLMs. We distinguish target models, tasks, motivations and datasets.
2 Preliminary
Threat model. We here consider a target LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$, with tokenizer $T$. The goal of the attacker $\mathcal{A}$ is now to leverage access to the LLM $\mathcal{M}$ to infer whether a specific piece of text has been part of its training data $\mathcal{D}$.
We note that MIAs against LLMs have been proposed across varying assumptions made for the attacker $\mathcal{A}$, from access to the model weights [55, 62], to only query access with model logits [106, 16, 68, 86, 69, 109, 101, 105, 108, 97, 73] or merely generated text [28, 45]. Some work also assumes access to additional non-member data from the same distribution as the target sample [101, 62, 97]. In this work, we focus on the evaluation of MIAs against LLMs, independent of the specific threat model applied.
Dataset Structure for LLM Training. We formalize the LLM training dataset $\mathcal{D}$ as a collection of $N$ documents: $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{N}\}$, where each document $d_{i}$ is composed of a sequence of tokens.
A document is an ordered sequence of tokens, $d_{i}=(t_{1},t_{2},...,t_{N_{d_{i}}})$, with $N_{d_{i}}$ the length of document $d_{i}$. We consider documents originating from a single, identifiable source which are coherent and semantically meaningful when considered in isolation. Examples include a news article, blogpost, book, academic paper, piece of code, etc. The document length can vary widely. For a single book this can be up to more than 100k tokens, while for social media posts this can be under $100$ tokens. The document length is a property of the document itself and is independent of whether the entire document fits into the LLM’s context window.
Given a document $d_{i}=(t_{1},t_{2},...,t_{N_{d_{i}}})$, a sequence $s_{ij}$ is a contiguous subsequence of tokens from $d_{i}$. Formally, for some starting index $a$ and ending index $b$, where $1\leq a\leq b\leq N_{d_{i}}$, a sequence $s_{ij}$ is defined as $s_{ij}=(t_{a},t_{a+1},...,t_{b})$. The set of all possible sequences from document $d_{i}$ is denoted as $\mathcal{S}_{i}$, where $\mathcal{S}_{i}=\{(t_{a},t_{a+1},...,t_{b})\mid 1\leq a\leq b\leq N_{d_{i}}\}$. The length of a sequence $s_{ij}$, denoted as $|s_{ij}|$, is defined as $b-a+1$. Sequences are not required to align with semantic boundaries in the document (e.g., sentences, paragraphs) and are defined by their start and end indices within the document. The sequence length can be up to the entire document ($a=1$ and $b=N_{d_{i}}$), although sequences are typically assumed to be shorter. The choice of sequence boundaries $(a,b)$ have been considered random (random excerpts from books) [86] or based on specific criteria (e.g. start of a document [27]).
LLM (pre)training. Tokenizer $T$ maps a piece of text $S_{ij}$ to a sequence of tokens $T(S_{ij})=s_{ij}=(t_{a},t_{a+1},\ldots,t_{b})$, where each token belongs to a vocabulary $\mathcal{V}$ of size $|\mathcal{V}|=V$.
Given a sequence of tokens (referred to as the context), auto-regressive language models predict the probability distribution for the next token over all tokens in vocabulary $\mathcal{V}$. This capability is acquired during training on the entire dataset $\mathcal{D}$, where the model minimizes the loss, generally measured using the cross-entropy function, for next-token prediction. For a model $\mathcal{M}$ with parameters $\theta$, we define the loss computed on sequence $s_{ij}$ as $\mathcal{L}_{\mathcal{M}}(s_{ij})=-\frac{1}{b-a+1}\sum_{k=a}^{b}\log\left(\mathcal{M}_{\theta}(t_{k}|t_{a},\ldots,t_{k-1})\right)$, where $\mathcal{M}_{\theta}(t_{k}|t_{a},\ldots,t_{k-1})$ represents the probability predicted by the model for the true token $t_{k}$ given the preceding tokens.
In this work, we exclusively consider models trained on causal language modeling objectives. This includes both base model versions — models trained on large quantities of unsupervised data — and versions that have been further fine-tuned on supervised task-specific data to unlock further abilities (instruction following, chat, specific tasks). In Table 1, we refer to these target models as pretrained and finetuned, respectively. Models trained with more complex alignment techniques such as reinforcement learning methods fall outside the scope of this work.
Sequence-level MIAs. The unit of interest in MIAs against LLMs has most commonly been a sequence. A sequence-level MIA thus aims to determine whether a given sequence $s_{t}$ has been used to train target LLM $\mathcal{M}$. Given an LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$ and a target sequence $s_{t}$, a sequence-level MIA is a scoring function $f_{\text{seq}}$ such that $f_{\text{seq}}(\mathcal{M},s_{t})$ outputs a continuous value representing the likelihood that $s_{t}$ was included in $\mathcal{D}$. A higher score indicates a higher likelihood of membership. To make a binary prediction $\hat{y}$, a threshold $\tau$ can be applied $\hat{y}_{\text{seq}}(\mathcal{M},s_{t})=1\text{ if }f_{\text{seq}}(\mathcal{M},s_{t})\geq\tau_{\text{seq}}\text{, and }0\text{ otherwise}$. The MIA performance can then be determined with the true label for membership $y$, $y_{\text{seq}}(\mathcal{M},s_{t})=1\text{ if }\exists d_{i}\in\mathcal{D},s_{t}\in d_{i}\text{, and }0\text{ otherwise}$.
MIAs are typically evaluated on the scoring function $f_{\text{seq}}$ using classification metrics that do not rely on a particular threshold, such as area under the receiver operating characteristic curve (AUC) or true positive rate (TPR) at low false positive rate (FPR) [12].
Most prior studies on MIAs against LLMs have considered sequences of a fixed, often truncated length significantly shorter than most documents. Carlini et al. [16] use sequences of $256$ tokens, while Shi et al. [86] and subsequent works [109, 101, 62, 105, 108, 97, 73, 45] consider random excerpts between $32$ and $256$ tokens and Meeus et al. [70] considers up to $100$ synthetically generated tokens. We thus primarily focus on sequence-level MIAs and discuss other units of interest in Sec. 5.2 and 6.
An MIA against an LLM thus comes down to designing the scoring function $f_{\text{seq}}(\mathcal{M},s_{t})$ which assesses the likelihood that a sequence $s_{t}$ was observed by model $\mathcal{M}$. Intuitively, model loss can be used as a naive scoring function (i.e. $f_{\text{seq}}(\mathcal{M},s_{t})=-\mathcal{L}_{\mathcal{M}}(s_{t})$[106], as model training minimizes entropy values for sequences included in the training data. Notably, individual sequences are not seen in isolation during training, as they are likely conditioned on preceding context. Especially for sequences smaller than documents (or data grouped during training), this might impact the effectiveness of using the model loss as a scoring function in sequence-level MIAs against LLMs compared to other ML models. Further MIAs against LLMs have proposed alternative scoring functions, for instance, by normalizing the model loss with respect to how ’normal’ a piece of text is [16], by contrasting the model loss with the loss of neighboring samples [68] or by only focusing on the lowest predicted probabilities [86].
3 Literature review
In Table 1, we review prior work studying MIAs against LLMs. We focus on the most recent work (’23-’24) and also include initial seminal work.
First, we distinguish between MIAs against finetuned and pretrained models, as defined in Sec. 2. We find that most recent MIAs are primarily targeting pretrained LLMs rather than models finetuned on smaller datasets.
Second, we distinguish between MIAs based on the unit of interest they consider. Most commonly, prior work has considered sequence-level MIAs as defined in Sec. 2. Some exceptions include scaling MIAs to the level of a document [86, 69] or dataset [67]. In the rest of this work we focus primarily on sequence-level MIAs, consistent with the majority of the literature, and will revisit document-level MIAs below.
Third, we report all studies developing new MIAs; i.e. designing novel scoring functions $f_{\text{seq}}(\mathcal{M},s_{t})$, which would more effectively capture the likelihood of membership [16, 106, 68, 86, 109, 109, 101, 62, 105, 108, 97, 73, 45, 55]. Studies not proposing new MIAs still evaluate MIAs on particular setups, e.g. to evaluate the effectiveness of privacy-preserving technologies [17, 79], detect the use of copyrighted content [70, 100] or draw connections between MIAs and PII leakage [65]. In Table 1 we show that there is a significant body of recent work focused on developing new MIAs.
Fourth, we distinguish the primary motivation of each MIA study as either privacy and security, copyright or dataset contamination. While initial work was primarily studying MIAs from a privacy and security perspective [15, 16, 71, 68, 65], we find that recent work increasingly considers novel questions associated with LLM pretraining, i.e. detecting the use of copyright-protected content or evaluation benchmarks.
Finally, across all works, we also classify the datasets which have been used to study MIAs. Importantly, we identify that most of the recent studies have relied on datasets collected post-hoc for MIA evaluation ().
4 Current evaluation methods are flawed
In this section, we examine the validity of using datasets collected post-hoc for MIA evaluation, as employed by most recent work ( in Table 1).
4.1 MIAs relying on post-hoc data collection
The evaluation of MIAs is typically instantiated as a randomized privacy game [104]. However, creating such a privacy game becomes challenging for LLMs, as the mere scale and cost of training prohibits the development of several LLMs (fixed target in [104]) or even one (fixed model in [104]) solely for the purpose of evaluating MIAs.
Recent work [86, 69] has developed strategies to create a set of likely members and likely non-members using information about training data sources and the LLM release date. We refer to such data collection strategies as post-hoc, where an attacker aims to infer membership for a released LLM, without access to the randomized training data distribution. In both works, members are collected by sampling documents from the often publicly available datasets widely used for LLM training (e.g. Wikipedia articles [86], books in the public domain from Project Gutenberg [^3] https://www.gutenberg.org/, ArXiv papers [^4] https://arxiv.org/ [69]). Non-members are collected as documents from the same data sources but released after the LLM’s training data cutoff date. Using this collection technique for evaluation, Shi et al. [86] propose a novel MIA scoring function based on the lowest predicted probabilities returned by the LLM. Authors achieved a membership AUC of $0.74$ for model LLaMA-30B [94] on the WikiMIA dataset, and an AUC of $0.88$ for GPT-3 [10] on copyright-protected books (BookMIA). Meeus et al. [69] train and evaluate a custom membership classifier against OpenLLaMA [32] and achieved MIA AUC of $0.86$ and $0.68$ for books and academic papers, respectively.
Such post-hoc approaches to collect datasets for MIA evaluation are not strictly randomized, raising the question whether they sufficiently approximate randomness in a privacy game or if biases may be introduced during the data collection process. Indeed, doubts have recently been raised on whether this construction of members and non-members constitutes a distribution shift due to the temporal gap in data collection [27, 67]. Even if efforts to mitigate any distribution shift are employed (e.g. Meeus et al. [69] ensured that the original publication dates of member and non-member books were similarly distributed), post-hoc data collection leaves the possibility of a distribution shift. If this is the case, the membership classifier might learn the temporal gap in language rather than membership, potentially questioning the validity of MIA results obtained.
4.2 Quantifying a potential distribution shift
Duan et al. [27] was the first to flag that the previously reported high MIA AUC might be due to the distribution shift between members and non-members. They observe a significant change in $7$-gram overlap between members and non-members, and also show how the MIA AUC increases when the temporal gap between members and non-members increases. Shortly afterwards, Maini et al. [67] also flags that post-hoc data collection introduces an important confounder, making members and non-members not IID.
We here propose to train and evaluate a bag of words classifier to distinguish members from non-members. We argue that such classifier can be used to measure what performance for this binary prediction task can be reached purely based on the textual data without taking the target model into account. If such baseline is better than a random guess, this would be indicative of a shift in language.
We use a random forest classifier ($500$ trees with maximum depth of $2$ and minimum samples per leaf of $10$) with the count of words ($1$-grams) appearing in at least $5\%$ of the training documents as features. For each dataset, we train the classifier on $80\%$ of all documents, evaluating it on the remaining $20\%$ (averaged over $5$ runs, with balanced membership).
We instantiate the bag of words classifier on multiple datasets collected post-hoc and used to evaluate MIAs against LLMs. We consider all (available) datasets: the WikiMIA and BookMIA datasets as introduced by Shi et al. [86], the books from Project Gutenberg and academic papers from ArXiv used by Meeus et al. [69], ArXiv papers from ArXivMIA [62] and StackExchange samples from StackMIA [105].
Dataset AUC Predictive words (sample)
WikiMIA (32) [86] $0.988\pm.004$2014,2023,2022}
WikiMIA (128) [86] $0.994\pm.003$2014,2023,2022, was}
BookMIA [86] $0.937\pm.004$didn,wasn,upon,hadn}
Gutenberg (full) [69] $0.997\pm.002$gutenberg,ebook,online,http}
Gutenberg (mid) [69] $0.889\pm.021$colored,n****,mediæval}
ArXiv [69] $0.690\pm.024$mathcal,2007,2010,prove}
ArXivMIA [62] $0.596\pm.037$training,mathbb,github}
StackMIAsub [105] $0.612\pm.004$magnetic,people,surface}
Table 2: AUC (mean±$\pm$±standard deviation) for the bag of words classifier across datasets collected post-hoc.
Table 2 shows that the bag of words classifier can distinguish members from non-members significantly better than a random guess (AUC of $0.5$). For some datasets, the classifier reaches near-perfect AUC, i.e. $0.99$ for the WikiMIA dataset, $0.94$ for the BookMIA dataset and $0.97$ for the books from Project Gutenberg (full), strikingly exceeding the MIA AUC reported by attacks utilizing the corresponding target model [86, 69]. For other datasets, the performance is relatively lower, yet remains significantly better (AUC of $\geq 0.6$) than a random guess baseline. This means that, all datasets relying on post-hoc data collection which have been released, suffer from a significant and often strong distribution shift. Moreover, the simplicity of the bag of words classifier suggests that it likely provides a lower bound on classification performance. More advanced classifiers could potentially detect other, subtler distribution shifts.
To further understand the effectiveness of the classifier and the underlying biases in the collected data, we examine the most predictive words for each dataset.
For WikiMIA [86], most predictive words correspond to years. As the dataset is constructed using text from Wikipedia ‘events‘, entries almost always contain dates. Data collected for different time periods can thus be perfectly distinguished by the bag of words classifier using primarily the years. Similarly for ArXiv papers used by Meeus et al. [69], years (in citations) carry the most predictive value. For ArXivMIA [62], the bias in years gets eliminated (they use abstracts, which rarely contain dates). Yet, we recover an AUC of $0.60$ with the most predictive words suggesting a change in topics over time (increased use of ’training’ and ’github’ in recent work).
Books, on the other hand, likely contain less dates. Meeus et al. [69] even attempted to control for a potential distribution shift by ensuring publication year to be similarly distributed between 1850 and 1910 for members and non-members. Yet, in this case, the AUC is near-perfect with the most predictive words relating to template text or formatting. We find that the beginning and end of the books from Project Gutenberg include certain templates (e.g. ‘This ebook was transcribed‘), which have changed over time, allowing the classifier to distinguish members from non-members. Similarly for BookMIA [86], certain formatting styles have changed between members and non-members (e.g. ‘did not‘ rather than ‘didn’t‘).
The distribution shift can be very subtle. We attempt to mitigate the distribution shift due to the template text by excluding the first and last 10% of the books (mid). Table 2 shows, however, that the AUC of the bag of words classifier does decrease from $0.97$, yet remains as high as $0.89$. When it comes to the most predictive words, we find that books, while originally published in the same time period but released online at a later stage, contain less old formatting and notably, less politically incorrect words such as the n-word.
Our results show that a simple bag-of-words classifier can consistently distinguish between members and non-members across all observed datasets. The analysis of the predictive words reveals that these biases can be subtle, and hard to mitigate. This implies that a fundamental assumption in membership inference, i.e. the members and non-members are IID, does not hold. The post-hoc collection of members and non-members thus fails to provide a proper, randomized privacy game to evaluate MIAs. We are hence unable to exclusively attribute the high MIA AUC reported by studies relying on post-hoc data collection [69, 86, 109, 101, 62, 105, 108, 97, 73, 45] to memorization of the LLM.
4.3 Perpetuating flaws in subsequent research
Since the initial work relying on data collected post-hoc [86, 69], we identify at least 10 novel attack methodologies evaluated using datasets collected in a post-hoc manner [83, 109, 101, 62, 28, 105, 108, 45, 97], the latest of which proposed (shortly before submission) in September ’24 [97]. This work either uses datasets released by Shi et al. [86] or newly created datasets of members and non-members in a similar, post-hoc manner: ArXivMIA [62], StackMIA [105], Dolma-Book [108] and Wikipedia Cutoff [73].
With the results of our bag of words classifier (Table 2), we find that all (accessible) datasets collected post-hoc suffer from distribution shift between members and non-members, which renders any comparison of MIA methodologies unreliable. Indeed, if one MIA method outperforms the other, it is impossible to determine whether one MIA is better at distinguishing which data points are memorized by the target LLM or if it is simply better at identifying the distribution shift.
The problem also extends beyond the development of new MIAs, with numerous subsequent work using similar datasets for evaluating the effectiveness of machine unlearning [103] or watermarking techniques [48, 100] as a privacy-preserving technology [79]. The use of this flawed MIA evaluation setup in broader contexts is equally concerning. It may lead to false conclusions about the effectiveness of privacy-preserving technologies, which could have far-reaching implications for LLM development and deployment.
5 Randomization for MIAs against LLMs
Before diving into the potential solutions for proper MIA evaluation, we examine the inherent challenges of creating a randomized setup for MIA evaluation.
5.1 Why is it different than for MIAs against ML models?
MIAs have been extensively applied to various domains of ML, including image classification [88, 14, 84, 96, 60] and natural language processing [85, 36, 93, 89], all evaluated in randomized privacy games. Ye et al. [104] distinguishes between a membership inference game involving a fixed record and one involving a fixed model. In the fixed record scenario, multiple ML models are trained with the target record either included or excluded, commonly referred to as shadow modeling [63, 64, 91, 41, 12, 82, 106]. For instance, Shokri et al.[88] used up to $100$ shadow models for image classification models, while Song et al.[90] trained 10 shadow models to evaluate privacy in smaller language models (RNNs). However, replicating this approach for LLMs is impractical due to the extremely high costs of pretraining. For example, training Llama-3.1-405B on 15 trillion tokens required 30.84 million GPU hours  [3]. Such resource demands make privacy games relying on multiple models infeasible.
As a result, privacy games to evaluate MIAs against LLMs likely rely on the fixed model scenario, in which a single ML model is trained with some randomly sampled target records from the same distribution, either included or excluded. While this setup has been widely examined in previous MIAs against ML [88, 18, 76, 84, 98], their practical implementation for LLMs is not without its limitations, which we will address below.
5.2 Unit of interest to infer membership
In traditional MIAs against ML models, the unit of interest to infer membership is naturally defined (e.g. single record, image). For text, most prior work considered sequences of a fixed, often truncated length as the unit of interest. For instance, Carlini et al. [16] use sequences of $256$ tokens, while Shi et al. [86] and subsequent work [109, 101, 62, 105, 108, 97, 73, 45] consider random excerpts between $32$ and $256$ tokens. When injecting randomized sequences, prior work has considered sequences up to $80$ [100] or $100$ [70] tokens.
In some cases, however, the motivation of applying MIAs against LLMs comes down to inferring membership of larger pieces of text, from the level of a document [69, 86] to an entire dataset [67]. For instance, in the context of copyright, the task of interest is inferring whether the entire content (e.g. news articles, book) was used rather than certain excerpts. This has motivated prior work [69, 86, 67] to instantiate document- or dataset-level membership inference, aiming to meaningfully aggregate membership signal to larger pieces of text.
A document-level MIA aims to determine whether a given document $d_{t}$ has been used to train target LLM $\mathcal{M}$. In Sec. 2, we formalize a document as an ordered sequence of tokens $d_{i}=(t_{1},t_{2},...,t_{N_{d_{i}}})$, of length $N_{d_{i}}$, which originates from a single, identifiable source and is coherent and semantically meaningful when considered in isolation. Given an LLM $\mathcal{M}$ trained on dataset $\mathcal{D}$ and a target document $d_{t}$, a document-level MIA is a scoring function $f_{\text{doc}}(\mathcal{M},d_{t})$ such that $\hat{y}_{\text{doc}}(\mathcal{M},d_{t})=1\text{ if }f_{\text{doc}}(\mathcal{M},d_{t})\geq\tau_{\text{doc}}\text{, and }0\text{ otherwise}$. The true membership label is $y_{\text{doc}}(\mathcal{M},d_{t})=1\text{ if }d_{i}\in\mathcal{D}\text{, and }0\text{ otherwise}$. For simplicity, we argue that the task of dataset inference as proposed by Maini et al. [67] is a generalization of document-level MIAs.
Document-level MIAs differ from sequence-level MIAs against LLMs, as aggregating meaningful signal across significantly more tokens might be challenging. Multiple ways have been proposed to compute $f_{\text{doc}}(\mathcal{M},d_{t})$. Meeus et al. [69] iterates through the entire document $d_{t}$, queries the LLM on truncated chunks, and aggregates all token-level probabilities to multiple document-level features as input for a membership classifier. Shi et al. [86] proposes to sample $k$ random excerpts, i.e. sequences $s_{tj}$, from document $d_{t}$ to compute the sequence-level scoring function $f_{\text{seq}}(\mathcal{M},s_{tj})$ for all $j=1,\ldots k$. The document-level scoring function is computed as an average binary prediction $y_{\text{doc}}(\mathcal{M},d_{t})=\frac{1}{k}\sum_{j=1}^{k}\hat{y}_{\text{seq}}(\mathcal{M},s_{tj})$, where $\tau_{\text{seq}}$ is determined to maximize accuracy on the sequence-level MIA. Lastly, Maini et al. [67] aggregates values of the sequence-level scoring function to a statistical test of the document level as $y_{\text{doc}}(\mathcal{M},d_{t})$.
5.3 Overlap between members and non-members
LLMs require massive datasets for pre-training, often collected from diverse internet sources (e.g. 15 trillion tokens for LLaMA-3 [3]). Since text is composed of sequences of tokens from a finite vocabulary, it tends to be lower-dimensional and more fragmented compared to data types like images. Given the sheer scale of these datasets, there is likely significant duplication of certain sequences. Moreover, text from the internet likely naturally contains significant duplication, with excerpts from e.g. books or news articles frequently appearing across various sources.
This inherent overlap of text is a well-known issue which has been heavily studied in the context of properly evaluating newly released LLMs. For instance, in GPT-3 [10], authors found a significant overlap between the training data and established evaluation benchmarks. To evaluate how the model generalizes to unseen data, they aggressively deduplicate ($13$-gram overlaps) between the training data and benchmarks. This methodology (decontamination) has since been adopted and further refined by subsequent model developers [99, 26, 95]. Even these ‘string matching’ methods however remain insufficient to address near-duplicates, semantic duplicates or translations [102].
The overlap also poses challenges for MIAs. Indeed, non-member data that is partially seen during training may act as quasi-members, complicating the inference process. Moreover, recent research suggests that near-duplicates in training fuels memorization. Shilov et al. [87] demonstrate that the inclusion of near-duplicates in the training dataset of an LLM enhances the distinguishability of a sequence in MIAs. Further, it is well established that deduplication, either literal [52] or semantic [1], improves LLM performance, suggesting that near-duplicates contribute to redundant memorization.
Recently, Duan et al. [27] applied common techniques used for decontamination on a randomized split of train and test data from the Pile [31]. These splits have been used to train and validate some open-source models [7, 8], which recent research [55, 67, 27] suggests to offer a randomized privacy game to evaluate MIAs against LLMs. Duan et al [27] finds that these splits, just as encountered for evaluation benchmarks, can indeed have substantial overlap, making it harder for MIAs to succeed. They report an average of 32.5% $7$-gram overlap between members and non-members for Wikipedia articles from the Pile [31], even reaching up to 76.9% overlap for data collected from Github.
6 Towards proper MIA evaluation
Aware of the importance of, and challenges associated with, creating a randomized setup to evaluate MIAs against LLMs, we here review and discuss potential solutions proposed in the literature.
6.1 Randomized test split
MIMIR [27] MIMIR [27]
Subset The Pile [2] (13_0.8) (7_0.2)
Wikipedia $0.497\pm.017$ $0.516\pm.040$ $0.580\pm.030$
Pile CC $0.551\pm.026$ $0.525\pm.028$ $0.539\pm.038$
PubMed $0.481\pm.010$ $0.508\pm.024$ $0.783\pm.017$
ArXiv $0.502\pm.017$ $0.501\pm.039$ $0.730\pm.013$
DM Math $0.503\pm.028$ $0.499\pm.025$ $0.799\pm.063$
HackerNews $0.488\pm.020$ $0.497\pm.021$
Github $0.507\pm.014$ $0.679\pm.028$ $0.860\pm.035$
Table 3: AUC (mean±$\pm$±standard deviation) for the bag of words classifier applied to subsets of The Pile, either directly sampled from its uncopryighted version [2] or from MIMIR [27] (which applied further deduplication).
The privacy game for a fixed model [104] only requires one training run, and could be naturally instantiated alongside regular training, making it feasible even for expensive LLMs.
Some pretrained LLMs make their entire training dataset open-source, including a randomized and controlled split of ’test’ data. While this set of test documents is primarily intended for a fair evaluation of the model’s ability to generalize its language modeling capabilities to unseen language, such test sets might prove very valuable to study and evaluate MIAs [27, 55, 67]. The training and validation splits are for instance made available for the suite of Pythia models [7], GPT-NEO [8], the OLMo model family [34] and the CroissantLLM project [30]. Finally, some initiatives such as OpenLLaMA [32] and LLM360 [61] release the exact training data sequences used, making it possible to know whether particular samples have been seen by intermediate model checkpoints or not.
First, we confirm whether the distribution shift identified earlier for members and non-members collected post-hoc, is now eliminated in a real-world, randomized setup. We consider three datasets. First, we consider $1,000$ members and non-members directly sampled from the train and test splits of (uncopyrighted version of) The Pile [2] (used to train and validate Pythia [7] and GPT-NEO [8]). Next, we consider both versions of the MIMIR dataset (13_0.8 and 7_0.2) [27]. As discussed in Sec. 5.3, Duan et al. [27] identified a substantial overlap between members and non-members in the randomized split from The Pile, making the notion of membership poorly defined and making it harder for MIAs to succeed. To mitigate this issue, they apply a deduplication strategy on the non-member data, removing non-members with more than 80% overlap in $13$-grams (13_0.8) and more aggressively, samples with more than 20% overlap in $7$-grams (7_0.2). Authors release the result as the MIMIR dataset. We train and evaluate a bag of words classifier on all three datasets and report the results in Table 3.
First, we confirm that, for the data directly sampled from the train and test split across all $7$ subsets from the Pile (left column), the bag of words classifier barely performs better than a random guess baseline; consistent with what we expect from a randomized split.
For the MIMIR dataset [27], however, the results (right columns) are more nuanced. For the moderate level of deduplication (13_0.8) for almost all data subsets, we find the AUC to remain close to a random guess baseline. This means that the distribution shift remains mitigated, while removing non-members with substantial overlap in the training dataset. For data collected from Github, however, the AUC of our bag of words classifier now corresponds to $0.68$, suggesting that there is a meaningful difference between members and non-members. From the analysis in Duan et al. [27], this suggests that the overlap between non-members and members in Github is so substantial that the non-members remaining after deduplication are mostly outliers, which initiates a new distribution shift. For the more aggressive deduplication (7_0.2), we find similarly large AUC for almost all subsets. This suggests that deduplicating too aggressively post-hoc might distort the exact properties of a randomized setup that are desired to evaluate MIAs. Future research is needed to understand the level of deduplication which can be applied post-hoc to mitigate the effect of the overlap issue, while not introducing any distribution shift.
Apart from the Github dataset split and the aggressive deduplication (7_0.2), the bag of words AUC suggests that the MIMIR dataset allows for a randomized setup to evaluate MIAs against LLMs. We instantiate a range of MIAs from the literature on the MIMIR dataset [27] (13_0.8, and not considering the Github subset) using Pythia-dedup-6.9B [7]. Table 4 summarizes the MIA AUC for this setup. We confirm, across a range of novel MIA methods proposed, the attack performance remains only barely better than random. Importantly, also the attacks proposed most recently [109, 101, 97, 105, 108, 45], which claim to out-perform prior methods in flawed evaluation setups, do not perform any better than a random guess in this clean setup.
We also instantiate document-level MIAs on a randomized test split of entire ArXiv documents from the Pile [2]. We consider both sequence-level MIAs from the literature, scaled to the document-level as proposed by Shi et al. [86] and formalized in Sec. 5.2 as well as the document-level MIA proposed by Meeus et al. [69]. Also for document-level MIAs, we find the MIA AUC to barely perform any better than random. We provide the results in Table 8 with the detailed experimental setup in (Appendix B).
Subset of The Pile [31]
MIA Wikipedia Pile CC PubMed Central ArXiv DM Math HackerNews
Bag of words $0.516\pm.040$ $0.525\pm.028$ $0.481\pm.010$ $0.502\pm.017$ $0.503\pm.028$ $0.488\pm.020$
Loss [106] $0.515\pm.013$ $0.512\pm.012$ $0.504\pm.013$ $0.525\pm.013$ $0.484\pm.013$ $0.517\pm.013$
Zlib [16] $0.516\pm.013$ $0.512\pm.013$ $0.506\pm.013$ $0.524\pm.013$ $0.486\pm.013$ $0.517\pm.013$
Lower [16] $0.518\pm.013$ $0.519\pm.013$ $0.513\pm.012$ $0.518\pm.012$ $0.493\pm.013$ $0.507\pm.012$
Ratio-LLaMA-2 [16] $0.529\pm.013$ $0.516\pm.013$ $0.524\pm.013$ $0.537\pm.013$ $0.503\pm.013$ $0.535\pm.013$
Neighborhood [68] $0.501\pm.013$ $0.511\pm.014$ $0.496\pm.013$ $0.530\pm.013$ $0.511\pm.013$ $0.484\pm.013$
Min-KProb [86] $0.515\pm.013$ $0.519\pm.012$ $0.510\pm.013$ $0.527\pm.013$ $0.492\pm.013$ $0.530\pm.013$
Min-K%++ [109] $0.510\pm.013$ $0.516\pm.013$ $0.506\pm.013$ $0.520\pm.012$ $0.491\pm.013$ $0.529\pm.013$
PAC [105] $0.482\pm.013$ $0.476\pm.013$ $0.479\pm.013$ $0.477\pm.013$ $0.5$013 $0.485\pm.013$
SURP [108] $0.511\pm.013$ $0.507\pm.013$ $0.504\pm.013$ $0.508\pm.013$ $0.492\pm.013$ $0.522\pm.013$
SaMIA [45] $0.468\pm.014$ $0.521\pm.012$ $0.518\pm.013$ $0.495\pm.013$ $0.493\pm.013$ $0.539\pm.013$
SaMIA*zlib [45] $0.507\pm.012$ $0.514\pm.012$ $0.527\pm.013$ $0.491\pm.013$ $0.524\pm.013$ $0.539\pm.013$
Table 4: Benchmarking MIAs on the MIMIR dataset [27] and Pythia-dedup-6.9B [7]. Sequences are sampled from the train (members) and test (non-members) split of The Pile [31]. MIA AUC computed using 500 members and 500 non-members.
Limitations. While held-out test splits generally provide a randomized setup for privacy games, the substantial overlap between members and non-members is a core limitation. Post-hoc deduplication could potentially mitigate the overlap, yet determining the appropriate level of deduplication is difficult to evaluate and might distort the randomization. Its effectiveness likely varies depending on the dataset type, as demonstrated in the case of the GitHub subset.
The practical significance of the overlap issue is a nuanced, perhaps philosophical, question. To rigorously examine what an LLM memorizes and how effectively MIAs can distinguish between members and non-members, one might prefer a setup with minimal overlap. Yet, in real-world applications, overlap is an inherent feature of text. To accurately infer whether a specific piece of text, such as a copyright-protected book or a medical record, was used to train an LLM, it is likely crucial to consider any overlap with other member data. For example, well-known quotes from a book might be widespread online (for which copyright might not apply), while generic elements of a medical record, even if present in the training data, may not pose a privacy risk. Whether MIAs -to e.g. study privacy risk in practice- should be evaluated on datasets with an overlap thus remains an open question.
6.2 Injection of randomized unique sequences
$n_{\text{rep}}$
MIA 10 100 1000
Bag of words $0.543\pm 0.37$ $0.543\pm 0.37$ $0.543\pm 0.37$
Loss [106] $0.509\pm.019$ $0.538\pm.018$ $0.592\pm.019$
Zlib [16] $0.514\pm.019$ $0.537\pm.019$ $0.569\pm.017$
Lower [16] $0.526\pm.018$ $0.548\pm.018$ $0.613\pm.018$
Ratio-LLaMA-2 [16] $0.542\pm.018$ $0.617\pm.018$ $0.704\pm.017$
Neighborhood [68] $0.488\pm.019$ $0.512\pm.019$ $0.566\pm.018$
Min-KProb [86] $0.509\pm.018$ $0.547\pm.018$ $0.626\pm.018$
Min-K%++ [109] $0.514\pm.018$ $0.531\pm.018$ $0.599\pm.018$
ReCaLL [101] $0.508\pm.019$ $0.538\pm.018$ $0.591\pm.018$
Probe (real) [62] $0.482\pm.018$ $0.460\pm.019$ $0.480\pm.018$
PAC [105] $0.512\pm.017$ $0.545\pm.018$ $0.587\pm.017$
SURP [108] $0.490\pm.018$ $0.550\pm.018$ $0.592\pm.019$
CON-ReCall [97] $0.504\pm.018$ $0.529\pm.018$ $0.477\pm.019$
SaMIA [45] $0.524\pm.018$ $0.582\pm.018$ $0.578\pm.018$
SaMIA*zlib [45] $0.578\pm.018$ $0.464\pm.019$ $0.538\pm.019$
Table 5: Benchmarking MIAs across copyright traps for CroissantLLM [30]: MIA AUC for $500$ members and non-members.
Instead of relying on a randomized test split of documents, prior work has proposed the injection of highly unique sequences, mitigating the overlap, to evaluate sequence-level MIAs.
For smaller language models, the controlled injection of specific sequences in the training data has been studied extensively. Henderson et al. [36] injects a variety of key-value pairs in a sequence-to-sequence dialogue model, while Carlini et al. [15] proposes the injection of canaries, unique sequences of random numbers, in the training set of LSTM models; both to measure how these sequences are memorized. Further work has, for instance, injected unique sequences to study memorization of federated learning [92] or of word embeddings [93].
More recently, Wei et al. [100] investigates how data watermarks, i.e. random sequences of tokens injected into the training dataset, are memorized by medium scale language models (70M parameters trained on 100M tokens). Further, Meeus et al. [70] injected canaries (which they call traps sequences) into the pretraining dataset of CroissantLLM [30], a 1.3B parameter model trained on 3T tokens. The canaries are synthetically generated, unique sequences, injected with varying sequence length, number of repetitions, and perplexity. As the canaries are synthetically generated (by sampling from a pretrained LLM), the same process allows generating new, unique canaries coming from the exact same distribution, yet not injected into the training dataset of CroissantLLM [30] (non-members).
While Meeus et al. [70] use this setup to study how the inclusion of canaries enables content owners to track whether their content has been used for LLM training, the setup also provides a randomized and controlled setup to evaluate MIAs against LLMs. We confirm this to be the case by fitting the bag-of-words classifier on copyright traps (using $500$ members and non-members), recovering an AUC of $0.543\pm 0.037$ for sequences of $100$ tokens. This confirms it to be a randomized setup for evaluating sequence-level MIAs.
Compared to the randomized setup available with a randomized test split (Sec. 6.1), the use of randomized unique sequences to evaluate MIAs might have two distinct benefits.
First, when the randomized sequences are highly unique, they could effectively eliminate the challenge of overlap between members and non-members. Meeus et al. [70] generate synthetic sequences of high perplexity, suggesting that these sequences would likely be more unique compared to the rest of the target model training data than a randomized test split.
Further, injected sequences can be repeated. As it is well established that repeated sequences are memorized more [13, 44], injecting with a varying number of repetitions would provide an MIA evaluation setup across multiple regimes of LLM memorization. This could also enable the comparison of different MIAs with (more) statistical significance.
We instantiate MIAs from the literature on the dataset of member and non-member canaries as made available by Meeus et al. [70] on the open-source CroissantLLM [30]. Table 5 summarizes the MIA AUC across a number of repetitions. For $1,000$ repetitions, we recover MIAs reaching a performance significantly better than a random guess, with Ratio-LLaMA-2 [16] (AUC of $0.70$) and Min-KProb [86] (AUC of $0.63$) the best performing methods.
Limitations. While generally a great randomized setup to develop and evaluate MIAs against LLMs, using the randomized unique sequences to evaluate MIAs also comes with its limitations.
First, the MIA performance might only represent a worst-case risk and be not representative of real-world data. Indeed, highly unique sequences mitigate any overlap between members and non-members, making it easier for MIAs to succeed, especially when repeated during training. Moreover, Meeus et al. [70] shows that sequences with larger perplexity are more vulnerable, and considers sequences with a perplexity higher than naturally occurring ones. Hence, evaluating MIAs on unique sequences might give a worst-case MIA risk compared to real-world data. While this might be beneficial for some use-cases (e.g. privacy auditing of methods with formal guarantees [63, 64, 91], detecting copyright-protected content [70, 100]), evaluating MIAs on (artificially) unique sequences might not be representative of risks in practice.
Second, the success of MIAs might depend on certain properties of the sequence. For instance, the Ratio MIA [16] relies on the use of a reference model, which should normalize the membership score by how ’complex’ a certain sequence is compared to other natural language. Surely, the extent to which this normalization is effective for an MIA against a target LLM depends on the choice of reference model, its training data and how well it generalizes to unseen data. Meeus et al. [70] use LLaMA-2 7B [95] to synthetically generate the trap sequences, which likely affects the success of certain reference models in detecting such traps. Also Duan et al. [16] finds "choosing the right reference model for a target LLM challenging and largely empirical". Moreover, we hypothesize that certain MIAs may vary in effectiveness based on sequence length (e.g. Min-KProb [86], which relies on a subset of token-level probabilities). Similarly, some MIAs (e.g. SMIA [73]) depend on the semantic meaning of a target sequence, which might be less effective for sequences with very high perplexity due to the lack of semantically meaningful embeddings in such cases. This suggests that using randomized unique sequences as an MIA evaluation method could influence the relative performance of MIA methodologies.
6.3 Finetuning
Thus far, we have worked under the assumption that training one (or more) models for the purpose of MIA evaluation is beyond reach for LLMs. While this is the case at the scale of modern LLM pretraining, it is generally feasible to run continued pretraining or to finetune LLMs on smaller datasets, and potentially using parameter-efficient finetuning methods [37, 39, 110, 23, 11, 58]. Researchers could thus finetune LLMs, with controlled injection of randomized members and non-members, to then evaluate MIAs on the finetuned models (fixed model) or even finetune multiple LLMs for a certain record (fixed record).
For instance, Mattern et al. [68] finetunes GPT-2 [81] to evaluate their MIA on a controlled set of members and non-members. Mireshghallah et al. [71] also finetunes GPT-2 models [81] and finds that finetuning the head of the model leads to better attack performance. Other work finetuned models to measure data extraction [107], or to apply MIAs for plagiarism detection [51] and studying PII leakage [65].
Leveraging finetuning for MIA evaluation has distinct benefits compared to previous approaches. First, similar to the injection of unique sequences with varying repetitions, finetuning could be done for varying memorization regimes (adjusting e.g. number of parameters updated during finetuning, dataset size, learning rate [71]). As such, this setup would be well suited to compare MIA performance across methods.
Further, from a practical perspective, it is very sensible to study MIAs in post-pretraining attack models. Finetuning (or continued pretraining) is often done on smaller, domain-specific, high quality datasets which are often more susceptible to contain sensitive information. It is cost effective to start off from generalist model checkpoints (e.g. LLaMA-3 [3], Mistral-7B [43]) and train on a domain-specific, causal language modeling task for a few billions of tokens to specialize models. Examples include finetuning for legal domain adaptation [19], for multilingual abilities [4], and for medical domain adapation [49]. Moreover, annealing, or post-training techniques in which LLMs are trained on high-quality data splits after their initial pretraining phase, is now understood to be essential to boost model performance [38, 3]. Such high-quality data (news articles, textbooks, synthetic data) is often the category most likely to be copyright-protected.
Limitations. For some MIA use-cases, including detecting the use of copyright-protected content or evaluation benchmarks during pretraining, the evaluation of MIAs remains generally desired for pretrained models rather than for finetuned models. How MIA evaluation on finetuned models translates to larger models, datasets or training hyperparameters remains an open question [6].
6.4 Post-hoc control methods
Throughout this work, we have made the case for randomized setups to evaluate MIAs against LLMs. We identified as potential solutions: (i) randomized test split, (ii) injection of randomized unique sequences and (iii) finetuning. However, such randomized setups are generally unavailable for most LLMs, especially for the widely used recent LLMs, such as Mistral-7B [43], LLaMA-2/3 [95, 3] and GPT-4 [77].
Therefore, we examine the options to still apply MIAs on real-world LLMs for which a controlled setup is unavailable. This might be useful to compare the memorization of certain LLMs or to calibrate MIAs (e.g. determining the threshold $\tau_{\text{seq}}$) before making an actual inference.
Unit of
interest Dataset Model(s) Benefits Limitations
Sequence The randomized split of The Pile [31], with additional deduplication, as made available by Duan et al. [27] in the MIMIR dataset [^5] https://huggingface.co/datasets/iamgroot42/mimir. We recommend using the split where sequences with more than 80% of $13$-gram overlapping are not removed, not considering data from Github. Suite of Pythia [7] and GPT-NEO [8] models. Allows for randomized MIA evaluation. In this setup, MIAs are more challenging due to the inherent overlap between members and non-members This however represent a more realistic setup in general.
Sequence The randomized unique traps sequences as made available by Meeus et al. [70], across all number of repetitions $(10,100,1,000)$. CroissantLLM [30] Allows for randomized MIA evaluation in a realistic real-world model, limiting the overlap between members and non-members, and across different memorization regimes. MIA performance might not be realistic, and might be specific to sequence characteristics such as length, and perplexity.
Document Randomized set of ArXiv papers sampled from the train and test splits from The Pile [31], both available as complete documents as well as in sequences. Suite of Pythia [7] and GPT-NEO [8] models. Allows for randomized MIA evaluation. In this setup, MIAs are more challenging due to the inherent overlap between members and non-members This however represent a more realistic setup in general.
Document Set of ArXiv papers sampled close to the training data cut-off date for OpenLLaMA[32]. Suite of OpenLLaMA[32] models. Allows for MIA evaluation on data collected post-hoc with any distribution shift between members and non-members largely mitigated (RDD). Might still be subject to a distribution shift, and faces the overlap issue.
Table 6: Recommended benchmarks to evaluate new MIAs against LLMs proposed in the future, highlighting benefits and limitations.
6.4.1 Beating the bag of words baseline
A potential solution is to revisit the baseline used to infer binary membership. MIAs against machine learning (ML) models traditionally use a 50/50 split of membership for evaluation, resulting in a random guess AUC of $0.5$ [88, 12, 18, 57, 84, 96, 35, 53, 75, 82, 106]. Alternatively, Jayaraman et al. [42] argues that a balanced prior for membership is not realistic in practice and proposes to evaluate MIAs with skewed priors. More recently, Kazmi et al. [47] proposes privacy auditing of ML models using a baseline from a (model-less) classifier.
Hence, we could set the performance of the bag of words classifier as a prior. An effective MIA methodology should then outperform this baseline by leveraging access to the target model and exploiting its memorization of member target records.
This solution has significant limitations in practice. First, the performances of the bag of words classifier applied to the widely used datasets for post-hoc MIAs (see Table 2) are near-perfect. With such a significant distribution shift between members and non-members, it would become practically impossible for novel MIAs to surpass this baseline and might lead us to (incorrectly) conclude that the target model does not memorize. Second, such model-less baselines will always be a lower bound. In Sec. 4, we only used a simple bag-of-words classifier, since its performance was already substantial. Had this classifier performed closer to random guessing, a subtler distribution shift might still exist, potentially detectable by more sophisticated classifiers [111, 24]. Hence, unless the distribution shift is strongly mitigated, using a model-less prior is not a valid option.
6.4.2 Regression Discontinuity Design (RDD)
In the domain of causal inference, Regression Discontinuity Design (RDD) is a technique used to study the effect of a treatment when the assignment to the treatment has not been randomized but is instead based on a cutoff point. By sampling observations in a principled manner close to this cutoff, RDD enables the estimation of the treatment effect even in the absence of a randomized setup. RDD is for instance widely used in econometrics [50], where cutoff points might include income levels for welfare programs or geographical borders when studying different policy implications.
We here examine if the RDD technique can be applied to the task of membership inference for LLMs. Specifically, we will sample data samples published close to the cutoff date for the LLM training data, attempting to mitigate any distribution shift between members and non-members.
We instantiate this RDD setup on OpenLLaMA [32], a model for which the entire training dataset is known. As members, we consider papers from a month before the cutoff date for the training dataset of the target LLM (February ’23). For non-members, we consider papers released in March ’23 — so one month after the member documents.
First, we apply the same bag of words model to quantify any distribution shift still present in the RDD data. We find an AUC of $0.534\pm 0.027$ when distinguishing members from non-members, suggesting that by sampling documents close to the cutoff date, we eliminate a substantial part of the previously present distribution shift for papers as used by Meeus et al. [69] (AUC of $0.69$) or for paper abstracts as used in ArXivMIA [62] (AUC of $0.60$).
Table 8 (Appendix B) summarizes the AUC for both sequence- as well as document-level MIAs across the literature applied to this setup. We find that the performance of MIAs remains close to a random guess baseline, as also found in the case of a randomized test split 6.1.
While the RDD method we propose seems to offer a viable solution for the setup at hand (using OpenLLaMa [32] and ArXiv papers), its practical implementation may face limitations. First, the RDD setup assumes knowledge of the training data cut-off date (which is likely quite different from the model release date) and relies on the availability of sufficient data released close to this date. For ArXiv papers, more than $10,000$ new papers are added every month  [^6] https://arxiv.org/stats/monthly_submissions, allowing for ample data collection close to the cut-off date. While similar arguments can be made for other frequently updated data sources (e.g. news articles), the RDD approach could prove challenging for less frequently updated sources (e.g. crime novels). Second, the RDD method aims to significantly mitigate distribution shifts between members and non-members. The quantification of such biases faces similar issues as discussed above. Lastly, the RDD setup does not address the issue of potential overlap between members and non-members. This might be preferred if one wants to calibrate realistic MIA performance in practice, but maybe less so when benchmarking new methods.
6.4.3 Debiasing techniques
Recently, Eichler et al.[29] introduced two novel methods to mitigate bias between members and non-members collected post-hoc. The first method creates “non-biased” datasets, where members and non-members share similar n-gram distributions. The second one creates “non-classifiable” datasets, which resist baselines not taking the target model into account such as the bag of words classifier we here propose. While authors show these techniques to eliminate biases present in the original datasets collected post-hoc, this approach faces similar limitations as other post-hoc control methods. Fully eliminating known biases, not introducing new biases and accurately evaluating their presence, remains a practical challenge.
7 Conclusion
In this work, we identified that recent research on MIAs against LLMs often uses flawed evaluation setups. In the last year, research has proposed at least $10$ different MIA methodologies, whose performance has been impossible to fairly and appropriately compare and benchmark.
After carefully examining potential solutions, and their drawbacks, we identify a set of benchmarks to use when new MIAs are proposed. We summarize our findings in Table 6.
For sequence-level MIAs, we propose a combination of the MIMIR dataset [27] (with limited deduplication) and the randomized trap sequences provided by Meeus et al. [70]. MIMIR enables a randomized setup for MIA evaluation across different model sizes [7, 8] and, despite its overlap issues, serves as a realistic benchmark for MIAs in practice. The trap sequences allow for benchmarking new MIA methodologies across different levels of memorization regimes; however, the corresponding results may depend on the specific properties of these sequences.
For document-level MIAs, utilizing a randomized split of ArXiv documents from The Pile along with the relevant models [7, 8] offers a robust evaluation framework. Additionally, sampling ArXiv documents close to the training data cut-off date for an open-source model (RDD) might provide a valuable supplementary setup.
When resources allow, and if the motivation of developing MIAs is not strictly tied to pretraining, finetuning on randomized sets of sequences or documents would also be recommended to compare new MIAs.
Finally, we encourage future research on MIAs against LLMs to always include a model-less baseline and to remain mindful of the inherent challenges (e.g. overlap, scaling from finetuning to pretraining) before interpreting results.
Availability. The code and data required to replicate the results in this paper, as well as to benchmark new MIAs against LLMs are available on Github.
Appendix A Detailed MIA implementation
Table 7 elaborates on the implementation details used for the MIAs from the literature we consider. Note that some MIAs, i.e. Loss [106], Zlib [16] and Lower [16], do not depend on the choice of certain hyperparameters.
MIA Hyperparameters and implementation details.
Ratio-LLaMA-2 [16] We use LLaMA-2 7B [95] as a reference model.
Neighborhood [68] We follow the neighborhood generation procedure outlined in Section 2.2 of Mattern et al. [68] with RoBERTa [59] as an MLM, neighborhood size of $50$ samples, dropout probability $0.7$ and top-k = $10$.
Min-KProb [86] Following the best performing setting reported in Section 4.3 of Shi et al. [86], we use $k=20$.
Min-K%++ [109] In line with Min-KProb, we use $k=20$.
ReCaLL [101] We use a fixed prefix with $10$ shots drawn from an auxiliary non-member distribution.
[…]
PAC [105] Using the implementation provided by Ye et al. [105], we generate $10$ augmentations with alpha parameter of $0.3$. We the compute polarized distance using $k_{m}ax=0.05$ and $k_{m}in=0.3$ following the best performing results reported in the original paper.
SURP [108] Following the best performing setup reported in Section 5.3 of Zhang et al. [108], we use $k=40$ with an entropy threshold of $2.0$
CON-ReCall [97] We use $10$-shot prefixes for both member- and non-member-conditioned likelihoods, drawn randomly and independently for each target example. Non-member prefixes are drawn from the auxiliary non-member distribution. For member prefixes we assume partial access by the attacker, and draw $10$ samples from the training dataset excluding the target example.
SaMIA [45] Using the default setup in the implementation provided alongside Kaneko et al. [45], we generate $10$ candidates using top-k sampling with top-k=$50$, temperature of $1.0$, and maximum length of $1,024$. We take the prefix to be $50$% of the input sample.
Table 7: MIA Implementation Details Summary
Note that we do not consider the MIAs requiring additional non-member data, i.e. ReCaLL [101], Probe (real) [62] and CON-ReCall [97], for the results in Table 4 as we use all non-member data for evaluation. Generating additional, synthetic non-member data for the trap sequences in Table 5 was feasible by following code made available by Meeus et al. [70].
Finally, for the results in tables 5 and 4, we compute the AUC on $1,000$ bootstrapped subsets of members and non-members and report both the mean and standard deviation across all results [5].
Appendix B Results for document-level MIAs
Data collection. We collect two datasets to evaluate document-level MIAs on LLMs:
  1. 1.
    ArXiv papers from the Pile. We randomly sample $1,000$ documents from the train set (members) and $1,000$ documents from the test set (non-members) from (the uncopyrighted version of) the Pile [2], ensuring that the selected documents have at least $5,000$ words (any sequences of characters seperated by a white space). We then split each document in $25$ sequences of $200$ words, and consider as full documents the first $5,000$ words.
  2. 2.
    ArXiv (RDD). We instantiate the RDD setup for the suite of OpenLLaMA [32] models, for which the entire training dataset is known. As members, we consider papers from a month before the cutoff date for the training dataset of the target LLM i.e. $13,155$ papers published in February 2023. We collect this set of documents using Github repository of RedPajama-Data [20]. For non-members, we consider the $16,213$ papers released in March 2023 - so one month after the member documents. We collect this set of documents through the Amazon S3 bucket hosted by ArXiv maintainers  [^7] https://info.arxiv.org/help/bulk_data_s3.html. We also apply the same preprocessing, and only consider documents of at least $5,000$ words, truncated to $5,000$ words and split into $25$ sequences of $200$ words.
Models. For the ArXiv papers from the Pile we use Pythia-dedup-6.9B [7] as target model, while for the RDD setup we use OpenLLaMA-7B [32].
MIA methodology. We instantiate sequence- and document-level MIAs from the literature. We report both the AUC achieved for all sequences ($50,000$ in total) as well as the AUC when the sequence-level membership scores are scaled up to the document-level. For this, we consider the method proposed by Shi et al. [86] and described in Sec. 5.2. Namely, a membership inference signal is computed on the $k=25$ sequences $s_{tj}$ of $200$ words per document $d_{t}$. Next, an optimal threshold $\tau_{\text{seq}}$ is determined for the inference signal to maximize accuracy on the sequence-level MIA. By classifying the selected sequences using the optimal threshold, the average binary prediction is then computed for a document, and is used as document-level scoring function $y_{\text{doc}}(\mathcal{M},d_{t})$, or $y_{\text{doc}}(\mathcal{M},d_{t})=\frac{1}{k}\sum_{j=1}^{k}\hat{y}_{\text{seq}}(\mathcal{M},s_{tj})$. Again, we do not consider the MIAs requiring additional non-member data, i.e. ReCaLL [101], Probe (real) [62] and CON-ReCall [97], and leave out SaMIA [45] and Neighborhood [68] due to the large computational costs required for this number of sequences.
We also instantiate the document-level MIA as proposed by Meeus et al. [69]. We query the target LLM on sequences of length $2,048$ tokens and consider both their ratio and maximum normalization strategy using token frequency (RatioNormTF and MaxNormTF), across two feature aggregations (simple aggregate statistics AggFE and histogram (HistFE)).
In all document-level MIAs, we consider $1,600$ documents for training (i.e. determining the optimal threshold $\tau_{\text{seq}}$ in the method by Shi et al. [86], and to train the meta-classifier for the method by Meeus et al. [69]) and use the remaining $400$ documents for evaluation (averaged over $5$ splits). For the sequence-level MIAs, as above, we compute the AUC on $1,000$ bootstrapped subsets of members and non-members and report both the mean and standard deviation across all results [5].
ArXiv (The Pile [31]) ArXiv (RDD)
MIA Sequence Document Sequence Document
Loss [106] $0.522\pm.003$ $0.527\pm.034$ $0.497\pm.002$ $0.512\pm.020$
Zlib [16] $0.527\pm.003$ $0.541\pm.028$ $0.501\pm.003$ $0.499\pm.022$
Lower [16] $0.514\pm.003$ $0.522\pm.028$ $0.508\pm.003$ $0.507\pm.024$
Ratio-LLaMA-2 [16] $0.537\pm.003$ $0.554\pm.030$ $0.516\pm.003$ $0.517\pm.003$
Min-KProb [86] $0.524\pm.003$ $0.525\pm.031$ $0.501\pm.003$ $0.496\pm.009$
Min-K%++ [109] $0.519\pm.003$ $0.520\pm.040$ $0.499\pm.003$ $0.494\pm.024$
SURP [108] $0.521\pm.003$ $0.526\pm.038$ $0$003 $0.487\pm.025$
AggFE+RatioNormTF [69] $0.548\pm.017$ $0.508\pm.046$
AggFE+MaxNormTF [69] $0.485\pm.013$ $0.522\pm.055$
HistFE+RatioNormTF [69] $0.531\pm.009$ $0.515\pm.026$
HistFE+MaxNormTF [69] $0.513\pm.026$ $0.513\pm.040$
Table 8: Benchmarking sequence and document-level MIAs. Left: MIA AUC for ArXiv papers ($25$ sequences of $200$ words for both $1,000$ members and non-members) randomly sampled from train and test split from The Pile [2], attacking Pythia-dedup-6.9B [7]. Right: MIA AUC for ArXiv papers (same length) randomly sampled from the RDD setup, attacking OpenLLaMA-7B [32].